# Enhancing LLM-based Code Translation in Repository Context via Triple Knowledge-Augmented

### v2

@@ 4.3. RQ3: Impact of Retrieval Strategy
Before Self-debugging After Self-debugging
Compilation Error Data type Misinterpretation 89 54
@@
As shown in Table 3, increasing the number of target-language code samples from one to two leads to improvements across Compilation@1, Pass@1, CodeBLEU, and Matchast, with Compilation@1 increasing notably from 64.8% to 68.0%. This indicates that adding more target-language code samples initially has a positive effect on the LLM. However, when the number is increased to three, the metrics no longer show significant improvements, and some, such as Compilation@1, even decrease, suggesting that excessive target-language code samples can introduce noise. Furthermore, as shown in RQ2 (Section 4.2), target-language code samples primarily impact match-based metrics like CodeBLEU and Match Score. Therefore, K3Trans ultimately sets the number of target-language code samples to two, as this configuration achieves the best performance on Compilation@1, Pass@1, CodeBLEU, and Matchast.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: The optimal number of target-language code samples to provide in the repository-context code translation scenario is two.
4.4. RQ4: Impact of Self-evolution
To evaluate the impact of the self-evolution process on K3Trans, we design three knowledge base stages—0%, 50%, and 100%—to simulate different stages of self-evolution. Here, 0% corresponds to the baseline (LLM-based with Prompt Engineering), and 100% corresponds to K3Trans. To comprehensively assess the practicality of the self-evolution process in different translation scenarios and to account for the randomness in the translation order of the target functions, we construct 50% knowledge base as follows: Target-language Code Samples: We randomly select 50% projects from the collected projects and extract functions to construct the samples. Dependency usage examples: For each target function, we randomly retain 50% of the dependency usage examples available at the 100% stage. The remaining 50% dependencies have their dependency usage examples set to empty, representing no reference is available in current project’s context. Successful translation function pairs: For each target function, we randomly select 50% of the successful translation function pairs available at the 100% stage. These selected pairs are then used as candidates for the top-1 successful translation function pair retrieval.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: As the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
4.5. RQ5: Bad Case Study
Figure 4. Example of Bad Case
@@ 4.5. RQ5: Bad Case Study
To understand the limitation of K3Trans, we further manually analyze the bad cases. In particular, we separately analyze the proportions of compilation errors and functional implementation errors in bad cases both before and after self-debugging. Furthermore, we classify compilation errors into finer-grained categories based on the taxonomy defined in RustRepoTrans (Ou et al., 2024). This analysis helps us identify which errors persist despite the triple-knowledge enhancement, as well as which errors can be easily fixed through self-debugging and which cannot. By doing so, we can pinpoint areas where future work should focus on improving LLM capabilities to further enhance performance in repository-level context code translation.
Before self-debugging, compilation errors account for 79.3%, which is already significantly lower than the 94.8% reported in (Ou et al., 2024) evaluating Basic LLM on RustRepoTrans. This indicates that triple knowledge enhancement effectively reduces issues such as dependency misuse, syntax misinterpretation, and syntactical difference confusion, thereby improving compilation rates. As shown in Figure 5, the highest proportion of compilation error types is Function Differences Misinterpretation, accounting for 59.3%, with a total of 309 occurrences (detailed in Table 5). Through case analysis, we found that K3Trans sometimes fails to provide LLM with fully corrective knowledge, particularly when the dependencies in the source project do not exist in the target project. as shown in Figure 4, for the source code statement: W[0].copy(this), K3Trans fails to provide the correct knowledge in both Target-Language Code Sample and Dependency Usage Example. This is because the dependency copy in the source project does not have a direct counterpart in the target project but should instead be implemented by using the built-in function clone in the target language. Since copy is not implemented in the target project, K3Trans cannot retrieve a valid dependency usage example for it. Additionally, a higher textual similarity does not always guarantee a functionally equivalent match. As a result, in some cases, the retrieved code sample fails to provide the correct translation guidance. Therefore, even though Successful Translation Function Pair provides a correct example, the noise introduced by the other two types of knowledge misleads the LLM, ultimately leading to an incorrect translation.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 5: K3Trans fails to provide LLM with fully corrective knowledge in certain cases, such as when the dependencies in the source project do not exist in the target project. However, many of the existing shortcomings in K3Trans can be effectively addressed by self-debugging.
@@ 5. THREATS TO VALIDITY
Threats in generalization. The only target language of the benchmark is Rust and the only tested LLM is Claude-3.5, which might limit the generalization of K3Trans. The reason of choosing the best-performing LLM, Claude-3.5, is that we aim to explore the upper limit of LLM performance on the challenging task of repository-level context code translation. This allows us to identify the gaps between the current capabilities of LLMs and the requirements of real-world translation scenarios. Additionally, our approach is not limited to Rust and can be extended to any other language in the future. This is because there is no language-specific adaptation in any phase of K3Trans. Therefore, to apply K3Trans to other programming languages, one only needs to collect relevant code for the target language and ensure the ability to analyze code in that language to extract target code sample. Extending the target language and the tested LLM would be part of our future work.
@@ 6.1. Code Translation
Code translation is a vital tool in software development, enhancing interoperability and enabling code reuse. It is crucial for updating legacy systems and integrating components in different languages. Research shows that code translation improves productivity by reducing manual translation time and costs (Lu et al., 2021; Puri et al., 2021), and supports the maintenance of multilingual projects.
@@ 6.2. Retrieval-Augmented Generation
Traditional LLMs like GPT and BERT rely on their training knowledge to handle tasks, exhibiting strong text generation but facing challenges like outdated knowledge, lack of domain expertise, and ”hallucinations” in specialized tasks. RAG addresses these by integrating external information sources into the model’s process (Khandelwal et al., 2019; Min et al., 2020; Zhang et al., 2025b; Shi et al., 2022; Liu et al., 2024). Specifically, RAG adds a retriever to the generative model, retrieving relevant data from external sources to improve performance.
@@
However, RAG faces challenges. One is maintaining a retrieval database, as Lin et al. (Lin et al., 2024) suggest addressing domain-specificity and knowledge base heterogeneity through continuous pretraining and a unified knowledge format. Another challenge is ensuring the retriever quickly identifies relevant documents, as Sawarkar et al. (Sawarkar et al., 2024) demonstrate by using semantic search and blended query strategies to improve retrieval performance. Our work constructs a self-evolving translation knowledge base to leverage triple knowledge to continuously improving translation quality of LLMs.
@@ 7. CONCLUSION
In this work, we propose a novel LLM-based code translation technique K3Trans, which leverages triple knowledge augmentation to enhance LLM’s translation quality under repository context in real-world software development. The experiments show that K3Trans substantially outperforms the baseline adapted from previous work by 19.4%/40.2% relative improvement in pass@1 and 0.138 in CodeBLEU. Furthermore, the code generated by K3Trans is of higher quality, as indicated by the higher DSR@1 metric of 82.1% and Repairable Ratio of 44.6%, which suggests a greater proportion of fixable code It is important to note that the results also demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Moreover, as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
Acknowledgements.

### v3

@@ 4.5.2. Analysis of Failure Examples
Before
Self-debugging
After
Self-debugging
Compilation
Error
Data type Misinterpretation 89 54
@@
Figure 6. Failure Example From K3Trans
Before self-debugging, compilation errors account for 79.3%, which is already significantly lower than the 94.8% reported in (Ou et al., 2024) evaluating Basic LLM on RustRepoTrans. This indicates that triple knowledge enhancement effectively reduces issues such as dependency misuse, syntax misinterpretation, and syntactical difference confusion, thereby improving compilation rates. As shown in Figure 8, the highest proportion of compilation error types is Function Differences Misinterpretation, accounting for 59.3%, with a total of 309 occurrences (detailed in Table 9). Through case analysis, we found that K3Trans sometimes fails to provide LLM with fully corrective knowledge, particularly when the dependencies in the source project do not exist in the target project. as shown in Figure 6, for the statement under translation: W[0].copy(this), K3Trans fails to provide the correct knowledge in both Target-Language Code Sample and Dependency Usage Example. This is because the dependency copy in the source project does not have a direct counterpart in the target project but should instead be implemented by using the built-in function clone in the target language. Since copy is not implemented in the target project, K3Trans cannot retrieve a valid dependency usage example for it. Additionally, a higher textual similarity does not always guarantee a functionally equivalent match. As a result, in some cases, the retrieved code sample fails to provide the correct translation guidance. Therefore, even though Successful Translation Function Pair provides a correct example, the noise introduced by the other two types of knowledge misleads the LLM, ultimately leading to an incorrect translation.
Finding 5: K3Trans fails to provide LLM with fully corrective knowledge in certain cases, such as when the dependencies in the source project do not exist in the target project. However, many of the existing shortcomings in K3Trans can be effectively addressed by self-debugging.
@@ 5. Threats To Validity
Threats in generalization. The only target language of the benchmark is Rust, which might limit the generalization of K3Trans. However, our approach is not limited to Rust and can be extended to any other language in the future, as there is no language-specific adaptation in any phase of K3Trans. Therefore, to apply K3Trans to other programming languages, one only needs to collect relevant code for the target language and ensure the ability to analyze code in that language to extract target code sample. Extending the target language would be part of our future work.
@@ 6.1. Code Translation
Code translation is a vital tool in software development, enhancing interoperability and enabling code reuse. It is crucial for updating legacy systems and integrating components in different languages. Research shows that code translation improves productivity by reducing manual translation costs (Lu et al., 2021; Puri et al., 2021), and supports the maintenance of multilingual projects.
@@ 6.2. Retrieval-Augmented Generation
Traditional LLMs like GPT and BERT rely on their training knowledge to handle tasks, exhibiting strong text generation but facing challenges like outdated knowledge, lack of domain expertise, and "hallucinations" in specialized tasks. RAG addresses these by integrating external information sources into the model’s process (Khandelwal et al., 2019; Min et al., 2020). Specifically, RAG adds a retriever to the generative model, retrieving relevant data from external sources to improve performance.
@@
However, RAG faces challenges. One is maintaining a retrieval database, as Lin et al. (Lin et al., 2024) suggest addressing domain-specificity and knowledge base heterogeneity through continuous pretraining and a unified knowledge format. Another challenge is ensuring the retriever quickly identifies relevant documents, as Sawarkar et al. (Sawarkar et al., 2024) demonstrate by using semantic search and blended query strategies to improve retrieval performance. Our work constructs a self-evolving strategy to continuously expand the translation knowledge base to consistently improve translation quality of LLMs.
@@ 7. Conclusion
In this work, we propose a novel LLM-based code translation technique K3Trans, which leverages evolving triple knowledge augmentation to enhance LLM’s translation quality under repository context in real-world software development. The experiments show that K3Trans substantially outperforms the baseline adapted from previous work, achieving relative improvements of up to 135.9% on Pass@1 and 32.8% on CodeBLEU among studied LLMs. Furthermore, the code generated by K3Trans is of higher quality, as indicated by the higher DSR@1 and Repairable Ratio, which suggests a greater proportion of fixable code. It is important to note that the results also demonstrate that each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution. Moreover, as the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.